OpenStudy (usarmy3947):

Marcos added +12 + -11 and got the answer -1. What error did Marcos make?

OpenStudy (texaschic101):

12 + (-11) -- a positive times a negative is a negative 12 - 11 = 1

OpenStudy (usarmy3947):

what?

OpenStudy (anonymous):

A positive times a negative is a negative. Your problem then turns into 12-11 = 1

OpenStudy (usarmy3947):

He found the difference of the numbers instead of the sum. He divided the integers. He chose the incorrect sign for the sum. He multiplied the integers.

OpenStudy (usarmy3947):

idk theres no numbers only this

OpenStudy (anonymous):

--- those are the answer choices, lol.

OpenStudy (usarmy3947):

yes

OpenStudy (anonymous):

12- -11=23 (A) 12/-11 = -1 (B) 12*-11=-132 (D)

OpenStudy (usarmy3947):

so its B

OpenStudy (usarmy3947):

He divided the integers.

OpenStudy (anonymous):

He DID "choose" the incorrect sign for the sum but there's more to it than just choosing. Although 12/-11 IS -1.09090909091, It could be C

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

I'd say C.. but wait to see if Ganeshie8 replies

OpenStudy (usarmy3947):

in the mean time can u help me with this one? Use the rules for operations with integers to simplify the expression. +16 × -2 -32 -8 +8 +32

OpenStudy (usarmy3947):

@Jemurray3 @Mp123

OpenStudy (anonymous):

I'm right here, why would you tag others? s: Are you sure it's +6 * -2?

OpenStudy (anonymous):

or is it +6x(-2)?

OpenStudy (usarmy3947):

its +6 X -2

OpenStudy (anonymous):

if it's the second one, it's -32. Courtesy of wolframalpha.com